# Astrophysical Black holes: An Explanation for the Galaxy Quenching

### v1

@@ Abstract
We propose models of Astrophysical Black holes (ABHs) without event horizons (EHs), as a more viable explanation for the long-term quenching phenomenon in galaxies. At the same time, the short-term quenching is explained here in terms of an efficient feedback expected in the models of stellar-mass astrophysical black holes (StMABHs). We have calculated the radiative flux from the disk in a general spherically symmetric metric background and used it to contrast the distinctive features of the BHs and ABHs scenarios. The nature of the feedbacks arising from accretion onto a BH and an ABH in the ‘quasar’ and ‘radio’ modes are compared and some possible observational signatures of the StMABHs are pointed out.
@@ I Introduction
In this paragraph, we outline the basic picture set out in the present study. The long-term quenching of galaxies is primarily attributed to a dense, compact object situated at the galactic nucleus[41]. Conventional black holes possess event horizon, limiting matter accretion up to a boundary which is set at the “Innermost stable Circular Orbit” (ISCO) and the consequent upper bound on the generation of ultra-high-energy particles. To address this, we propose here models of “astrophysical black holes” (ABHs) that are devoid of an event horizon (EH), hence characterized by a huge amount of matter confined within an extremely compact region, a situation which permits the disk of accreting matter to extend all the way to the centre and yield extremely high energy densities and temperatures. This access to the ultra-strong gravity regime facilitates generation of extremely energetic particles, thus contributing to the long-term quenching of star formation in the host galaxy and its halo region. Additionally, we introduce models of “stellar-mass astrophysical black holes” (StMABHs) which can arise from gravitational collapse of massive stars once they run out of their internal nuclear fuel. These StMABHs, prior to getting enveloped by the event horizon, create high-density regions capable of emitting very high-energy particles over short time scales. Such a situation may lead to short-term episodes of (localised) intra-galactic quenching. Potential observational signatures associated with the StMABHs are also pointed out here. Section II provides a brief overview of the recent observations related to this phenomenon, made with the James Webb Space Telescope (JWST). Section III touches upon the role SMBHs play in the galaxy quenching process. Section IV presents some salient features of our proposed models involving astrophysical black holes (ABH), in the context of long- and short-term quenching. Our conclusions are summarised in Section V.
@@ IV.1 A plausible scenario for long-term quenching
Long-term quenching can be attained through a sustained vigorous emission of high-energy particles emanating from the central entity of the galaxy. Suppose the central entity is an astrophysical black hole, an ultra-compact object lacking an event horizon. In that case, the accretion disk will extend toward the centre, as opposed to its terminating at the ISCO (see above). As we shall argue, this scenario has the potential to offer a more plausible and realistic explanation for the long-term quenching, as it permits extremely high-energy particle emissions over a long period of time. In 2011 and 2014, two such models were introduced, each suggesting the formation of an ultra-dense region through the dynamical collapse of certain distinct types of matter, such as a perfect fluid and others. These Joshi-Malafarina-Narayan(JMN) models posit astrophysical black holes (ABHs) which lack an event horizon, thus allowing the accretion disk to extend essentially right up to the center[50, 51].
@@ IV.3 The Role of Radiative and Mechanical Efficiency
IV.4 Outflows from Naked Singularities
@@ IV.5 Accelerating Winds
IV.6 Some Observational Signatures
@@ V Summary and Conclusions
We have adopted the basic calculation of Silk and Rees to show that the ABHs will theoretically require about 17 times more mass than the BHs to expel all the material from their host galaxies. The reason lies in the $\sim 100\%$ radiative efficiency of the accretion disk around ABHs, as explained in Section C. However, we have also presented the general calculation of the radiation output in an accretion disk model and discussed that ABHs would have an increasing radiative efficiency in the formation phase of the ultra-dense, compact object, and it will be 100% once the singularity is formed in the infinite comoving time. The main difference between the quenching due to the BHs and ABHs can be understood through their behaviors in the ‘quasar’ and ‘radio’ modes. In the ‘quasar’ mode, when the accretion rate of an AGN is very high, the BHs are very efficient in producing the radiation-driven winds as the radiative efficiency of the BHs is itself very low (i.e.,   6%), hence the requirement of a much higher accretion rate in the case of BHs. However, an ABH is forming ultra-dense compact objects at the center, so its radiative efficiency is increasing as the ISCOs get closer and closer to the center. In fact, ABHs are efficient producers of both the radiation-driven and mechanically driven winds in the quasar mode. On the other hand, the ‘radio’ mode differs in the two cases, as the BHs produce mechanically driven winds and/or jets with greater efficiency, whereas the ABHs are more efficient in producing radiation-driven winds. Still, the radiative efficiency of the ABHs is reaching the maximum.

### v2

@@ Abstract
In light of increasing observational evidence supporting the existence of ultra-compact objects, we adopt the term astrophysical black hole to refer to any object having a huge mass confined within a sufficiently small region of spacetime. This terminology encompasses both the classical black hole solutions predicted by general relativity, as well as alternative compact objects that may not possess an event horizon. We propose models of Astrophysical Black holes (ABHs) without event horizons (EHs), as a more viable explanation for the long-term quenching phenomenon in galaxies. At the same time, the short-term quenching is explained here in terms of an efficient feedback expected in the models of stellar-mass astrophysical black holes (StMABHs). We have calculated the radiative flux from the disk in a general spherically symmetric metric background and used it to contrast the distinctive features of the BHs and ABHs scenarios. We demonstrate the relative ease of wind generation from the accretion disk surrounding an ABH without an event horizon, compared to a BH, and highlight the significant strength of these winds. The nature of the feedbacks arising from accretion onto a BH and an ABH in the ‘quasar’ and ‘radio’ modes are compared and some possible observational signatures of the StMABHs are pointed out.
@@ I Introduction
In this paragraph, we outline the basic picture set out in the present study. The long-term quenching of galaxies is primarily attributed to a dense, compact object situated at the galactic nucleus[41]. Conventional black holes possess event horizon, limiting matter accretion up to a boundary which is set at the “Innermost stable Circular Orbit” (ISCO) and the consequent upper bound on the generation of ultra-high-energy particles. The cosmic censorship conjecture[42, 43] remains a foundational yet unproven and not even properly formulated, assumption in black hole physics. While it holds in simplified models such as spherically symmetric dust collapse, a wide range of studies on more general collapse scenarios—including those involving inhomogeneous matter fields, pressure, and rotation—have shown that naked singularities can emerge as legitimate solutions to Einstein’s equations [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Despite multiple attempts, no rigorous mathematical proof exists for any version of cosmic censorship, and the conjecture’s vague formulation continues to hinder progress. In this light, the theoretical exclusion of naked singularities is no longer viable. Since the formation of such singularities is a generic outcome in many physically reasonable collapse models, the focus now necessarily shifts toward their observational signatures. In this work, we consider the astrophysical implications of naked singularities—viewed as ultra-compact, massive objects without event horizons—and explore their potential impact on galaxy evolution, particularly in comparison to traditional black holes. To address this, we propose here models of “astrophysical black holes” (ABHs) that are devoid of an event horizon (EH), hence characterized by a huge amount of matter confined within an extremely compact region, a situation which permits the disk of accreting matter to extend all the way to the centre and yield extremely high energy densities and temperatures. This access to the ultra-strong gravity regime facilitates generation of extremely energetic particles, thus contributing to the long-term quenching of star formation in the host galaxy and its halo region. Additionally, we introduce models of “stellar-mass astrophysical black holes” (StMABHs) which can arise from gravitational collapse of massive stars once they run out of their internal nuclear fuel. These StMABHs, prior to getting enveloped by the event horizon, create high-density regions capable of emitting very high-energy particles over short time scales. Such a situation may lead to short-term episodes of (localised) intra-galactic quenching. Potential observational signatures associated with the StMABHs are also pointed out here. Section II provides a brief overview of the recent observations related to this phenomenon, made with the James Webb Space Telescope (JWST). Section III touches upon the role SMBHs play in the galaxy quenching process. Section IV presents some salient features of our proposed models involving astrophysical black holes (ABH), in the context of long- and short-term quenching. Our conclusions are summarised in Section V.
@@ IV.1 A plausible scenario for long-term quenching
Long-term quenching can be attained through a sustained vigorous emission of high-energy particles emanating from the central entity of the galaxy. Suppose the central entity is an astrophysical black hole, an ultra-compact object lacking an event horizon. In that case, the accretion disk will extend toward the centre, as opposed to its terminating at the ISCO (see above). As we shall argue, this scenario has the potential to offer a more plausible and realistic explanation for the long-term quenching, as it permits extremely high-energy particle emissions over a long period of time. In 2011 and 2014, two such models were introduced, each suggesting the formation of an ultra-dense region through the dynamical collapse of certain distinct types of matter, such as a perfect fluid and others. These Joshi-Malafarina-Narayan(JMN) models posit naked singularities, which are the astrophysical black holes (ABHs) that lack an event horizon, thus allowing the accretion disk to extend essentially right up to the center[68, 69].
@@ IV.3 The Role of Radiative and Mechanical Efficiency
IV.4 Wind Generation
We can get the idea of the wind generation due to the effects of the radiation pressure through the mechanism discussed in [77]. The key assumptions of the mechanism are as follows: Particles are ejected from the surface of a geometrically thin disk around a central object of mass $M$. The gravitational force of the central object, the centrifugal force, and the radiative forces from the disk act on the particles. The gravitational effects of the disk gas are neglected. The radiation force, the radiation drag from the central object, and the radiation drag due to the disk radiation field and the very high velocity motion of the particles are also considered zero. The near disk approximation is used for simplicity, in which the particles are assumed to suffer from the radiation fields of the disk just below them. The motions of the particles are calculated using the Newtonian formalism. In the cylindrical coordinates $(r,\phi,z)$, where $z$ axis is along the axis of the rotation of the disk, the equations of motion for the particles are[77],
$\frac{dv_{r}}{dt}=\frac{v^{2}_{\phi}}{r}-\frac{GMr}{R^{3}},$ (14)
$\frac{1}{r}\frac{d}{dt}(rv_{\phi})=0,$ (15)
$\frac{dv_{z}}{dt}=-\frac{GMz}{R^{3}}+\frac{\sigma_{T}}{mc}F_{d}.$ (16)
Here, $R=\sqrt{r^{2}+z^{2}}$ and $F_{d}$ represents the radiation flux from the disk and can be expressed as,
$F_{d}=\sigma T^{4}_{d}=\frac{3GM\dot{M}}{8\pi r^{3}}\left(1-\sqrt{\frac{r_{ISCO}}{r}}\right),$ (17)
where, $\dot{M}$ represents the constant accretion rate, $r_{ISCO}$ is the inner most stable circular orbit and $T_{d}$ is the effective temperature of the disk. The Eqn.(14) represents the equation of motion in the radial direction, with the first term being the centrifugal force and the second being the gravitational force due to the central object, on the right-hand side. The Eqn.(15) is the angular momentum conservation equation. The Eqn.(16) is the equation of motion in the vertical direction, with the first term being the gravitational force due to the central object and the second being the radiation force due to the disk radiation, on the right-hand side. We can define the normalized disk luminosity as,
$\Gamma_{d}=\chi_{e}\frac{L_{d}}{L_{E}},$ (18)
with ionization rate $\chi_{e}=1$, Eddington luminosity $L_{E}=4\pi GcMm_{p}/\sigma_{T}$ (here $m_{p}$ is the proton mass), and the disk luminosity,
$L_{d}=\frac{GM\dot{M}}{2r_{ISCO}}.$ (19)
In terms of $\Gamma_{d}$, Eqn.(16) can be rewritten as,
$\frac{dv_{z}}{dt}=-\frac{GMz}{R^{3}}+\frac{3\Gamma_{d}GMr_{ISCO}}{r^{3}}\left(1-\sqrt{\frac{r_{ISCO}}{r}}\right).$ (20)
To understand wind generation, we need to examine the escape condition for the particles. This can be achieved using the net specific energy of the particles ejected from the disk. For the particles ejected from the disk with the initial conditions: $v_{r}=0$, $v_{\phi}=v_{Keplerian}=\sqrt{GM/r}$, $v_{z}=0$, and $r=r_{0}$ at $z=0$ this specific energy will be of the order of
$E=\frac{v^{2}_{\phi}}{2}-\frac{GM}{r_{0}}+\frac{3\Gamma_{d}GMr_{ISCO}}{r_{0}^{3}}\left(1-\sqrt{\frac{r_{ISCO}}{r_{0}}}\right)z.$ (21)
The right-hand side consists of the rotational energy in the first term, the gravitational potential energy in the second, and the work done by the radiation force of the disk. If we measure $r_{0}$ in terms of $2GM/c^{2}$ and velocity in terms of $c$, the Eqn.(21) will take the form,
$E=-\frac{1}{2r_{0}}+\frac{3}{2}\Gamma_{d}\frac{r_{ISCO}}{r_{0}^{3}}\left(1-\sqrt{\frac{r_{ISCO}}{r_{0}}}\right)z.$ (22)
The gravitational acceleration is largest when $z\sim r_{0}$, hence demanding this energy to be positive at $z\sim r_{0}$ gives the escape condition as,
$\Gamma_{d~{}escape}=\frac{1}{6}\frac{r_{0}}{r_{ISCO}}\frac{1}{\left(1-\sqrt{\frac{r_{ISCO}}{r_{0}}}\right)}.$ (23)
Figure 2: The shaded regions show the normalized luminosity $\Gamma_{d}$ for which the particles will escape from their respective positions $r_{0}$. (a)Escape condition for ABH without EH case (b)Escape condition for BH case.
In the units of $2GM/c^{2}$, the ISCO for the Schwarzschild black hole is located at $r_{0}=3$. However, the ISCO for the ABH without EH models discussed in the previous section would be very close to the central ultra-dense object. If we assume the ultra compact object to be of the size of the order of Planck length, then ISCO for a few solar mass ABH without EH will be at $r_{0}=10^{-38}$. This value is inversely proportional to the mass of the ultra-compact object. The behavior of the escape condition for BH and ABH without EH shows a similar trend of the required disk luminosity diverging as the position of the particle reaches the ISCO, as shown in the Figs.2(a) & 2(b). However, as the disk extends to the central ultra compact object in ABH without EH, the divergence of the normalized luminosity occurs in the ultra strong gravity region compared to the Schwarzschild black hole where this divergence occur at the ISCO, which is very far from the EH. It is evident from Eqns. (18) & (19), when the disk luminosity is low due to the low accretion rate, the wind generation from the BH is not possible. However, even with the very low accretion rate, the ABH without EH will be able to generate the winds due to their extremely small value of the ISCO and hence the large value of $\Gamma_{d}$. The Eqn.(18) shows that the $\Gamma_{d(ABH~{}without~{}EH)}\sim 3\times 10^{38}\Gamma_{b(BH)}$ for the same rate of accretion. Hence, the winds generated from the accretion disk around the ABH without EH will be much stronger compared to those generated from the disk around a BH. These calculations neglect general relativistic effects; however, they offer a rough estimate of the relative ease of wind generation from ABHs without an event horizon compared to BHs, and consequently, their role in the quenching of star formation.
IV.5 Outflows from Naked Singularities
@@ IV.6 Accelerating Winds
IV.7 Some Observational Signatures
@@ V Summary and Conclusions
We have adopted the basic calculation of Silk and Rees to show that the ABHs will theoretically require about 17 times more mass than the BHs to expel all the material from their host galaxies. The reason lies in the $\sim 100\%$ radiative efficiency of the accretion disk around ABHs, as explained in Section C. However, we have also presented the general calculation of the radiation output in an accretion disk model and discussed that ABHs would have an increasing radiative efficiency in the formation phase of the ultra-dense, compact object, and it will be 100% once the singularity is formed in the infinite comoving time. The main difference between the quenching due to the BHs and ABHs can be understood through their behaviors in the ‘quasar’ and ‘radio’ modes. In the ‘quasar’ mode, when the accretion rate of an AGN is very high, the BHs are very efficient in producing the radiation-driven winds as the radiative efficiency of the BHs is itself very low (i.e.,   6%), hence the requirement of a much higher accretion rate in the case of BHs. However, an ABH is forming ultra-dense compact objects at the center, so its radiative efficiency is increasing as the ISCOs get closer and closer to the center. In fact, ABHs are efficient producers of both the radiation-driven and mechanically driven winds in the quasar mode. On the other hand, the ‘radio’ mode differs in the two cases, as the BHs produce mechanically driven winds and/or jets with greater efficiency, whereas the ABHs are more efficient in producing radiation-driven winds. Still, the radiative efficiency of the ABHs is reaching the maximum. We have examined wind generation from the accretion disk around a Schwarzschild black hole (BH) and ABHs without an event horizon. Our analysis shows that while the escape conditions exhibit similar behavior in both cases, ABHs are more efficient at producing winds, particularly at low accretion rates, where BHs fail to do so. This efficiency enables ABHs to drive winds even with minimal accretion, contributing to the quenching of star formation.